# Make Your Guard Learn to Think: Defending Against Jailbreak Attacks with Safety Chain-of-Thought

### v1

@@ 3.3 Supervised Fine-Tuning
Our objective is to train the model to proactively assess the harmfulness of each input before generating responses. Training exclusively on refusal reasoning risks conditioning the model to reject all inputs indiscriminately. To counteract this, we implement supervised fine-tuning using our SCoT dataset, complemented by a dataset of benign samples to retain balanced decision-making capabilities.
Retain Dataset Construction. We construct a benign dataset $\mathcal{D}_{b}=\{(q)\}^{k}$, containing $k$ benign questions. The augmented answer is divided into two components: the safety reasoning process and the original output generated by our target model $\mathcal{M}$. This reasoning process is illustrated as follows:
#### Verify Request
I need to verify the dangerousness of the
users request to [summary] across the
following categories before generating the
response.
* harassment/threatening: No.
* hate/threatening: No.
* illicit/violent: No.
* self-harm/intent/instructions: No.
* sexual/minors: No.
* violence/graphic: No.
* misinformation: No.
* privacy/sensitive: No.
Therefore, the users request is safe, and
I can proceed to generate the response.
#### Response
[output]
@@
$=-\sum_{t=1}^{T}\log p(y_{i,t}\mid x_{i},y_{i,<t};\theta)$
$\quad-\lambda\sum_{t=1}^{T}\log p(y_{j,t}\mid x_{j},y_{j,<t};\theta)$ (2)
@@ 4.1 Experiment setup
Training Dataset. We employ the circuitbreaker dataset introduced by (zou2024circuitbreaker) as our base dataset $\mathcal{D}_{r}$ for further development. This dataset comprises 4,994 short harmful requests across 48 harmful topics. For the retain dataset $\mathcal{D}_{b}$, we utilize dolly-15k (DatabricksBlog2023DollyV2) to preserve the general capabilities of the models. The Dolly-15k dataset is an open-source instruction-following records with diverse categories such as brainstorming, classification, closed QA, generation, information extraction, open QA, and summarization.
@@
SCoT Construction. We employ GPT-4o-mini (openai2024gpt4technicalreport) to evolve both the base questions and the answers as detailed in Sections 3.2 and 3.1 respectively. For adapting the questions, we select demonstrations of different styles from sorrybench (xie2024sorrybenchsystematicallyevaluatinglarge) and conduct few-shot In-Context learning. Detailed construction process are provided in Appendix A.
@@
Training Base Models and Hyperparameters. We employ two open-source models specifically tuned for safety, including Llama3-8B-Instrct (llama3modelcard) and Mistral-7B-V0.2 (jiang2023mistral7b), without system prompts. Training is conducted using LoRA (hu2021loralowrankadaptationlarge). For the LoRA module, we specify a rank of 64, an $\alpha$ value of 128, a dropout rate of 0.1, and learned LoRA matrices for all attention matrices. In the supervised fine-tuning stage, we set $\lambda=1$ in Equation 1 and the training epoch to 2. The initial learning rate is set to $52e-$ 5 and the batch size to 2. Training is carried out on two Ada 6000 GPUs and takes approximately two hours to complete.
Baselines. We evaluate our SCoT by comparing it with three distinct baselines. The first includes the original Mistral and Llama-3 Instruct models, serving as the simplest comparison point. Next, we assess against the circuitbreaker approach (zou2024circuitbreaker), which mitigates harmful content by projecting harmful representations into a randomized space, effectively creating a protective “circuit” against harmful requests. Finally, we compare our results to an adversarially trained version of the Mistral model, known as R2D2 (mazeika2024harmbenchstandardizedevaluationframework). We obtain the baseline models from the Huggingface repository (wolf2020huggingfacestransformersstateoftheartnatural), as released by the original authors.
Table 1: Evaluation of LLM jailbreak attack success rates (ASR). The term “None” denotes verbatim harmful requests. Labels (A) and (J) correspond to the evaluation results from Advbench and Jailbreakbench, respectively. For GCG attack for baselinses, results are taken from the reported findings in the Circuitbreaker paper (zou2024circuitbreaker) and are indicated with a superscript *.
Llama3-8B Mistral-7B
ASR($\downarrow$) Base Model advTrain circuitbreaker SCoT Base Model advTrain circuitbreaker SCoT
None(A) 0.07 - 0.01 0.01 0.45 0.44 0.00 0.00
None(J) 0.07 - 0.00 0.00 0.46 0.41 0.00 0.00
slang 0.08 - 0.09 0.02 0.22 0.28 0.09 0.02
uncommon dialect 0.08 - 0.07 0.00 0.33 0.25 0.09 0.04
expert endorsement 0.11 - 0.12 0.00 0.09 0.01 0.02 0.01
evidence-based 0.05 - 0.17 0.00 0.09 0.02 0.06 0.00
role play 0.12 - 0.05 0.02 0.69 0.47 0.08 0.00
logical appeal 0.03 - 0.18 0.00 0.09 0.03 0.03 0.00
affirmative(A) 0.01 - 0.00 0.01 0.92 0.01 0.00 0.04
affirmative(J) 0.07 - 0.00 0.06 0.94 0.01 0.05 0.02
suppress refusal(A) 0.18 - 0.13 0.02 0.76 0.21 0.00 0.00
suppress refusal(J) 0.37 - 0.03 0.04 0.77 0.40 0.06 0.03
GCG 0.45* - 0.03* 0.08 0.89* 0.08* 0.11* 0.20
@@
Evaluation Metrics. To evaluate the output harmfulness, the conventional method is substring matching using predefined refusal prefixes. However, the circuitbreaker approach ceases output upon detecting harmful content instead of issuing refusal answers. Consequently, we employ the Llama3Guard classifier (dubey2024llama3herdmodels) to assess the Attack Success Rate (ASR), determining whether generated outputs contain harmful content. As for evaluating general capabilities, we use accuracy as the metric.
@@ 4.2 Jailbreak Evaluation
Setup. We summarize the jailbreak attacks used for our evaluations, organized by the strategy and datasets employed:
  • No attack: As a control, we use a non-jailbreaking approach that echoes each prompt verbatim, facilitating baseline comparisons. Evaluations are performed using JailbreakBench (chao2024jailbreakbenchopenrobustnessbenchmark) with 100 harmful questions and AdvBench (zou2023universal) with 500 harmful behaviors.
  • Linguistic manipulation: We assess our SCoT in defending against linguistic manipulations, such as slang and uncommon dialects. This evaluation employs the sorrybench dataset (xie2024sorrybenchsystematicallyevaluatinglarge), which includes 100 harmful questions for each category.
  • Contextual manipulation: The model’s resilience to distracted content is tested using tactics like role play, logical appeal, expert endorsement, and evidence-based persuasion from the sorrybench dataset. We further assess our SCoT’s capability to suppress refusal and encourage affirmative responses, applying these strategies to the JailbreakBench and AdvBench datasets (wei2024jailbroken).
  • Adaptive attack: We employ a fuzzing-based white-box adaptive attack, GCG (zou2023universal), running for 500 iterations with a batch size of 512. The fitness function evaluates the loss of mutations. The implementation is based on the Harmbench (mazeika2024harmbenchstandardizedevaluationframework) framework.
Results. Table 1 presents the jailbreak attack success rates (ASR) of our SCoT compared to baselines across various attack scenarios. Our SCoT demonstrates a significant improvement in defense, reducing ASR to near zero for direct harmful requests and effectively countering both linguistic and contextual manipulation attacks. In contrast, models such as advTrain, based on Mistral-7B (mazeika2024harmbenchstandardizedevaluationframework), and Circuitbreaker (zou2024circuitbreaker), though trained on a diverse set of harmful prompts, exhibit limited adaptability to novel attack scenarios.
However, SCoT does show a higher ASR than Circuitbreaker in the GCG attack when using the Mistral-7B-based model. Upon further analysis of the compromised samples, we found that all failures fall within the misinformation category. We hypothesize that this may be due to the inherent weaker reasoning ability of the Mistral-7B-based model in specific domains. Overall, SCoT’s robust mechanism for detecting malicious intent allows it to generalize effectively to unseen attacks, including those targeting refusal suppression and eliciting affirmative responses. These findings underscore the superior generalization capability of SCoT across a wide range of jailbreak scenarios.
@@ 4.3 Potential Compromise in General Capabilities
Setup. To evaluate potential compromises in general capabilities due to security enhancements, we employ two significant benchmarks: MMLU (hendrycks2021measuringmassivemultitasklanguage) and GSM8K dataset (cobbe2021training). MMLU includes multiple-choice questions across 57 tasks such as elementary mathematics, US history, and computer science. GSM8K is designed to assess model performance on complex problem-solving tasks typical of graduate-level exams.
Table 2: Evaluation of LLM general capabilities using our SCoT and three baseline models. Performance is measured by accuracy, utilizing GPT-4o-mini as the answer cleansing model.
Capability($\uparrow$) llama3-8b mistral-v0.2-7b
base model advTrain circuitbreaker scot base model advTrain circuitbreaker scot
gsm8k 85.2% - 79.6% 83.4% 49.2% 0.0% 48.2% 47.0%
mmlu 72.8% - 65.0% 71.2% 56.2% 0.0% 53.6% 54.8%
Table 3: Ablation Study Results: Impact of Retain Dataset and Question Variants on Model Performance.
Llama3-8B Mistral-7B
base w/o retain dataset w/o question variants ours base w/o retain dataset w/o question variants ours
Capability($\uparrow$) gsm8k 0.85 0.33 0.85 0.83 0.49 0.45 0.47 0.47
mmlu 0.73 0.37 0.67 0.71 0.56 0.34 0.55 0.55
Robustness($\downarrow$) None(A) 0.07 0.01 0.01 0.01 0.45 0.08 0.07 0.00
None(J) 0.07 0.00 0.01 0.00 0.46 0.05 0.07 0.00
slang 0.08 0.02 0.00 0.02 0.22 0.03 0.03 0.02
uncommon dialect 0.08 0.01 0.02 0.00 0.33 0.03 0.02 0.04
expert endorsement 0.11 0.01 0.00 0.00 0.09 0.00 0.03 0.01
evidence-based 0.05 0.00 0.01 0.00 0.09 0.00 0.04 0.00
role play 0.12 0.02 0.03 0.02 0.69 0.01 0.04 0.00
logical appeal 0.03 0.00 0.00 0.00 0.09 0.01 0.02 0.00
affirmative(A) 0.01 0.02 0.23 0.01 0.92 0.03 0.02 0.04
affirmative(J) 0.07 0.01 0.18 0.06 0.94 0.04 0.07 0.02
suppress refusal(A) 0.18 0.04 0.02 0.02 0.76 0.07 0.07 0.00
suppress refusal(J) 0.37 0.04 0.01 0.04 0.77 0.21 0.14 0.03
Results. Table 2 reveals the trade-offs encountered when bolstering safety defenses. It shows the accuracy metrics across the GSM8K and MMLU datasets. Our SCoT demonstrates minimal performance loss relative to the base model, in contrast to Circuitbreaker, which shows more degradation. This drop in accuracy is likely due to an over-refusal response to questions containing terms that could violate safety protocols.

### v2

@@ 3.3 Supervised Fine-Tuning
Our objective is to train the model to proactively assess the harmfulness of each input before generating responses. Training exclusively on refusal reasoning risks conditioning the model to reject all inputs indiscriminately. To counteract this, we implement supervised fine-tuning using our evolved dataset, complemented by a dataset of benign samples to retain balanced decision-making capabilities.
Retain Dataset Construction. We construct a benign dataset $\mathcal{D}_{b}=\{(q)\}^{k}$, containing $k$ benign questions. Mimicing Fig 4, the augmented answers for benign samples are divided into two components: the safety reasoning process and the original output generated by our target model $\mathcal{M}$. An example of the reasoning process is shown in Appendix B.1.
@@
$=-\sum_{t=1}^{T}\log p(y_{i,t}\mid x_{i},y_{i,<t};\theta)-\lambda\sum_{t=1}^{T}\log p(y_{j,t}\mid x_{j},y_{j,<t};\theta)$ (2)
@@ 4.1 Experiment setup
Training Dataset. We employ the circuitbreaker dataset introduced by [63] as the base dataset $\mathcal{D}_{r}$ for further development. This dataset comprises 4,994 short harmful requests across 48 harmful topics. For the retain dataset $\mathcal{D}_{b}$, we utilize dolly-15k [10] to preserve the general capabilities of the models. The Dolly-15k dataset is an open-source instruction-following records with diverse categories such as brainstorming, classification, closed QA, generation, information extraction, open QA, and summarization.
@@
SCoT Construction. We employ GPT-4o-mini [32] to evolve both the base questions and the answers as detailed in Sections 3.2 and 3.1 respectively. For evolving the questions, we select demonstrations of different styles from sorrybench [48] and conduct few-shot In-Context learning (ICL). Detailed construction process is provided in Appendix A.
@@
Training Base Models and Hyperparameters. We employ two open-source models specifically tuned for safety, including Llama-3.1-8B-Instruct [1] and Mistral-7B-Instruct-v0.2 [21], without system prompts. Training is conducted using LoRA [19]. For the LoRA module, we specify a rank of 64, an $\alpha$ value of 64, a dropout rate of 0.1, and learned LoRA matrices for all attention matrices. In the supervised fine-tuning stage, we set $\lambda=1$ in Equation 1 and the training epoch to 3. The initial learning rate is set to 2e52e-5. Training is carried out on two Ada 6000 GPUs and takes approximately three hours to complete.
Baselines. We evaluate our SCoT by comparing it with four distinct baselines. First, we include the original Mistral-7B-Instruct-v0.2 and Llama-3.1-8B-Instruct models as simple, unaligned baselines for comparison. Second, we evaluate a prompting-based baseline that directly appends safety-related instructions to the user input [49]. While such prompting may appear unnatural under certain attack settings, it helps assess the necessity of SFT process. To account for variability in prompt design, we manually construct three prompt types: explicit step-by-step instructions (P1), structured output format instructions (P2), and conversational-style guidance (P3). The details of the prompts are provided in Appendix C. Third, we compare against the RRapproach [63], which suppresses harmful responses by projecting harmful activations into a randomized subspace, acting as a form of neural “circuit breaker”. Finally, we include R2D2 [30], an adversarially trained variant of the Mistral-7B-Instruct-v0.2 model, designed to enhance robustness against harmful queries. All baseline models are obtained from the Hugging Face repository [47].
Table 1: Evaluation of LLM Jailbreak ASR. “Verbatim” denotes direct, unaltered harmful requests and serves as a control baseline. Attack methods are sourced from Sorrybench[48] and Jailbroken[45]; detailed descriptions can be found in Appendix D.1. For adaptive attacks on baseline models, results marked with a superscript * are drawn from the reported findings in [63]. Bold indicates the best performance.
ASR()($\downarrow$) Mistral-7B-Instruct-v0.2 Llama-3.1-8B-Instruct
Original $R$ R2D2 P1 P2 P3 SCoT Original $R$ P1 P2 P3 SCoT
Verbatim 0.45 0.00 0.23 0.01 0.01 0.00 0.00 0.04 0.01 0.01 0.05 0.04 0.00
slang 0.22 0.06 0.25 0.02 0.03 0.00 0.00 0.04 0.09 0.00 0.04 0.01 0.00
uncommon dialect 0.33 0.10 0.27 0.02 0.04 0.05 0.00 0.02 0.07 0.01 0.05 0.00 0.00
translate-fr 0.40 0.21 0.37 0.09 0.05 0.01 0.01 0.06 0.06 0.02 0.05 0.03 0.00
translate-ml 0.55 0.43 0.19 0.06 0.03 0.02 0.00 0.38 0.09 0.05 0.17 0.07 0.03
translate-ta 0.26 0.21 0.08 0.06 0.05 0.02 0.00 0.34 0.09 0.09 0.22 0.08 0.02
translate-mr 0.29 0.11 0.11 0.07 0.10 0.02 0.01 0.22 0.09 0.01 0.23 0.11 0.05
translate-zh-cn 0.35 0.24 0.17 0.06 0.05 0.02 0.03 0.14 0.13 0.01 0.09 0.03 0.00
misspellings 0.43 0.15 0.24 0.06 0.05 0.05 0.00 0.07 0.11 0.01 0.08 0.02 0.00
disemvowel 0.29 0.22 0.02 0.07 0.00 0.01 0.00 0.18 0.10 0.03 0.02 0.04 0.00
leetspeak 0.45 0.39 0.00 0.10 0.01 0.00 0.00 0.06 0.07 0.00 0.01 0.01 0.00
expert endorsement 0.13 0.05 0.14 0.11 0.08 0.08 0.02 0.09 0.12 0.10 0.03 0.01 0.00
evidence-based 0.09 0.04 0.08 0.07 0.05 0.05 0.00 0.09 0.17 0.07 0.05 0.00 0.00
role play 0.72 0.05 0.63 0.08 0.09 0.17 0.02 0.14 0.05 0.04 0.10 0.09 0.00
logical appeal 0.15 0.02 0.08 0.10 0.06 0.07 0.00 0.05 0.18 0.08 0.05 0.03 0.00
misrepresentation 0.12 0.05 0.11 0.09 0.08 0.04 0.00 0.05 0.35 0.12 0.05 0.02 0.00
authority_endorsement 0.15 0.07 0.15 0.02 0.11 0.12 0.01 0.14 0.17 0.08 0.05 0.00 0.00
technical terms 0.53 0.06 0.47 0.12 0.07 0.04 0.00 0.04 0.29 0.06 0.11 0.04 0.01
prefix_injection 0.92 0.06 0.16 0.02 0.00 0.04 0.00 0.05 0.12 0.02 0.02 0.00 0.00
refusal_suppresion 0.76 0.11 0.15 0.05 0.04 0.01 0.00 0.11 0.11 0.00 0.04 0.00 0.00
style_injection_short 0.90 0.18 0.25 0.03 0.05 0.05 0.00 0.14 0.14 0.00 0.04 0.04 0.01
style_injection_json 0.95 0.11 0.36 0.03 0.00 0.02 0.01 0.07 0.28 0.00 0.05 0.00 0.00
distractors 0.27 0.04 0.23 0.01 0.00 0.00 0.00 0.00 0.02 0.00 0.05 0.02 0.00
poems 0.19 0.25 0.11 0.01 0.00 0.01 0.00 0.05 0.00 0.00 0.07 0.01 0.00
GCG 0.89* 0.11* 0.08* - - - 0.38 0.45* 0.03* - - - 0.00
AutoDAN 0.93* 0* 0* - - - 0.00 0* 0* - - - 0.00
PAIR 0.70* 0.23* 0.60* - - - 0.00 0.19* 0.08* - - - 0.00
@@
Evaluation Metrics. To evaluate the output harmfulness, we employ the Llama3Guard classifier [28] to assess the Attack Success Rate (ASR), determining whether generated outputs contain harmful content. As for evaluating general capabilities, we use accuracy as the metric.
@@ 4.2 Jailbreak Evaluation
Setup. We summarize the jailbreak attacks here and provide full details of them in Appendix D.1.
Verbatim  To establish a baseline for comparison, we employ a control method that directly echoes each prompt verbatim. This non-jailbreaking approach is evaluated against 500 harmful behaviors from the AdvBench dataset [64].
Linguistic manipulation  We evaluate SCoT’s robustness against a wide range of linguistic manipulations. This includes slang, uncommon dialects, and cross-lingual translations, sourced from SorryBench [48]. We also incorporate misspelling, disemvowel, and leetspeak attacks, following the setup in Jailbroken [45]. Cipher-based encodings such as Base64 and ASCII are excluded, as our experiments show that the models generally fail to decode these formats correctly and instead generate meaningless outputs according to our manual inspection.
Contextual manipulationSCoT’s resilience to contextual manipulation is tested using tactics from the SorryBench dataset [48], such as role-playing, logical appeals, expert endorsements, evidence-based persuasion, misrepresentation, authority endorsements, and the use of technical terms. Furthermore, we evaluate SCoT’s ability to counter strategies like refusal suppression, prefix injection, style injection, and distraction techniques. These latter strategies, drawn from Jailbroken [45], are applied to prompts from the AdvBench dataset.
Adaptive attack  We employ: GCG[64], a fuzzing-based white-box attack; AutoDAN[27], an evolutionary algorithm-based method; and PAIR [6], an iterative refinement approach utilizing Large Language Models (LLMs). The implementation is based on the Harmbench [30] framework.
Results. Table 1 reports the ASR of SCoT compared to baseline methods across 26 attack scenarios. SCoT consistently achieves near-zero ASR on direct harmful prompts and demonstrates strong robustness against linguistic, contextual, and adaptive attacks. In contrast, models like R2D2 [30] and RR [63], despite being trained on diverse harmful inputs, struggle to generalize to OOD attacks.
Prompting-based defenses perform comparably in some settings but suffer from inconsistent effectiveness and require careful prompt engineering. SCoT outperforms all prompting variants across the board, highlighting the limitations of direct prompting and the necessity of supervised fine-tuning.
One exception occurs under the GCG attack on the Mistral-based model, where SCoT shows slightly higher ASR than RR. Upon manual inspection, these failures primarily fall under the misinformation and illicit/violent categories. We hypothesize that this may be due to the inherent weaker reasoning ability of the Mistral-7B-based model in specific domains.
@@ 4.3 Potential Compromise in General Capabilities
Setup. To evaluate potential compromises in general capabilities due to security enhancements, we employ two significant benchmarks: mmlu [18] and gsm8k dataset [9]. mmlu includes multiple-choice questions across 57 tasks such as elementary mathematics, US history, and computer science. gsm8k is designed to assess model performance on complex problem-solving tasks typical of graduate-level exams.
Table 2: Evaluation of LLM general capabilities using our SCoT and three baseline models. Performance is measured by accuracy (%), utilizing GPT-4o-mini as the answer cleansing model.
Capability($\uparrow$) Mistral-7B-Instruct-v0.2 Llama-3.1-8b-instruct
Original $R$ R2D2 SCoT Original $R$ SCoT
gsm8k 49.5 45.8 38.6 47.0 85.7 79.6 85.3
mmlu 56.3 55.6 55.3 53.6 69.7 63.0 67.2
Table 3: Ablation Study Results: Impact of Retain Dataset (R), Question Variants (V) and CoT Reasoning (CoT) on Model Performance.
Mistral-7B-Instruct-v0.2 Llama-3.1-8b-Instruct
Original w/o R w/o V w/o CoT SCoT Original w/o R w/o V w/o CoT SCoT
Capability($\uparrow$) gsm8k 0.50 0.00 0.47 0.47 0.47 0.86 0.00 0.82 0.85 0.85
mmlu 0.56 0.00 0.57 0.55 0.54 0.73 0.00 0.69 0.65 0.66
Robustness($\downarrow$) Verbatim 0.45 0.00 0.00 0.03 0.00 0.04 0.00 0.00 0.02 0.00
slang 0.22 0.00 0.00 0.01 0.00 0.04 0.00 0.00 0.02 0.00
uncommon dialect 0.33 0.00 0.00 0.02 0.00 0.02 0.00 0.00 0.06 0.00
translate-fr 0.40 0.00 0.02 0.12 0.01 0.06 0.00 0.01 0.03 0.00
translate-ml 0.55 0.00 0.02 0.11 0.00 0.38 0.00 0.11 0.25 0.03
translate-ta 0.26 0.00 0.01 0.01 0.00 0.34 0.00 0.06 0.23 0.02
translate-mr 0.29 0.00 0.02 0.02 0.01 0.22 0.00 0.01 0.12 0.05
translate-zh-cn 0.35 0.00 0.01 0.11 0.03 0.14 0.00 0.01 0.08 0.00
misspellings 0.43 0.00 0.01 0.08 0.00 0.07 0.00 0.00 0.04 0.00
disemvowel 0.29 0.00 0.01 0.06 0.00 0.18 0.00 0.00 0.01 0.00
leetspeak 0.45 0.00 0.00 0.03 0.00 0.06 0.00 0.01 0.01 0.00
expert endorsement 0.13 0.00 0.05 0.01 0.02 0.09 0.00 0.01 0.01 0.00
evidence-based 0.09 0.00 0.01 0.01 0.00 0.09 0.00 0.01 0.01 0.00
role play 0.72 0.00 0.01 0.01 0.02 0.14 0.00 0.03 0.01 0.00
logical appeal 0.15 0.00 0.02 0.00 0.00 0.05 0.00 0.03 0.02 0.00
misrepresentation 0.12 0.00 0.05 0.00 0.00 0.05 0.00 0.04 0.02 0.00
authority_endorsement 0.15 0.00 0.05 0.02 0.01 0.14 0.00 0.03 0.00 0.00
technical terms 0.53 0.00 0.02 0.02 0.00 0.04 0.00 0.01 0.04 0.01
prefix_injection 0.92 0.00 0.01 0.04 0.00 0.05 0.00 0.01 0.01 0.00
refusal_suppresion 0.76 0.00 0.00 0.09 0.00 0.11 0.00 0.00 0.04 0.00
style_injection_short 0.90 0.00 0.02 0.13 0.00 0.14 0.00 0.01 0.01 0.01
style_injection_json 0.95 0.01 0.00 0.05 0.01 0.07 0.00 0.02 0.03 0.00
distractors 0.27 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00
poems 0.19 0.00 0.00 0.04 0.00 0.05 0.00 0.00 0.01 0.00
GCG 0.89* 0.00 0.50 0.88 0.38 0.45* 0.00 0.08 0.00 0.00
AutoDAN 0.93* 0.04 0.00 0.08 0.00 0* 0.00 0.13 0.00 0.00
PAIR 0.70* 0.00 0.08 0.08 0.00 0.19* 0.00 0.08 0.00 0.00
Results. Table 2 reveals the trade-offs encountered when bolstering safety defenses. It shows the accuracy metrics across the gsm8k and mmlu datasets. Our SCoT demonstrates minimal performance loss relative to the base model in most cases, in contrast to RR, which shows more degradation.